OpenStudy (anonymous):

Factor by grouping: 6k^2 + 5k - 4 So far I know that the sum = 4 and the product = 24, therefore the integers are 8 and -3 So now the equation is: 6k^2 + 8k -3k - 4

hero (hero):

Factor the first two terms: 6k^2 and 8k Factor the last two terms -3k - 4 Then notice what is common to both factorizations and factor that out as well.

hero (hero):

For example: x^2 + 5x + 6 Notice that: 3 × 2 = 6 3 + 2 = 5 Therefore: x^2 + 5x + 6 = x^2 + 3x + 2x + 6 = x(x + 3) + 2(x + 3) = (x + 3)(x + 2)

OpenStudy (anonymous):

Oh alright that makes sense, I'm just having trouble making both factorizations equal: \[6k ^{2} + 8k - 3k - 4\] \[-3k (-k + 1) + 4 (-2k + 1)\] Shouldn't they both be (-k + 1) or (-2k + 1)? I think I'm just doing something wrong

hero (hero):

Yes, you can only factor out what is common to both terms. What is common to 6k^2 and 8k ? To figure that out you would need to know what is common to 6 and 8. Then what is common to k^2 and k. To figure out what is common to 6 and 8, perform an appropriate factorization of each number. Notice that 6 = 2(3) 8 = 2(4) k^2 = k(k) k = k(1) From this, you can conclude that 2k is common to 6k^2 and 8k. When you factor that out, you get 6k^2 + 8k^2 = 2k(3k + 4)

OpenStudy (anonymous):

Alright thank you so much